OpenStudy (anonymous):

please help! hour and a half till the question is due. not sure what i'm doing wrong here. thanks (1 pt) Suppose F⃗ (x,y)=−xyi⃗ +y^2j⃗ . (a) Find a vector parametric equation for the parabola y=x^2 from the origin to the point (4,16) using t as a parameter. r⃗ (t)= (b) Find the line integral of F⃗ along the parabola y=x^2 from the origin to (4,16).

OpenStudy (anonymous):

the first part i got as x=t and y=t^2. or <t,t^2>

OpenStudy (anonymous):

tried \[\int\limits_{0}^{4}(-t)(t^2)+(t^2)^2 = 704/5\] for part b but it's not right.

OpenStudy (sidsiddhartha):

thers a small mistake that is u forget to multiply r(t)=<1,2t)\[F=-xy \hat{i}+y^2 \hat{j}\\x=t,y=t^2\\F_n=(-t^3 \hat{i}+t^4 \hat{j})\\r=<t,t^2>\\dr=<1,2t>=\hat{i}+2t \hat{j}\]

OpenStudy (sidsiddhartha):

so \[line~Integral=\int\limits_{C}^{}F_n.dr=\int\limits_{}^{}(-t^3 \hat{i}+t^4 \hat{j})*(\hat{i}+2t \hat{j})\]

OpenStudy (sidsiddhartha):

you missed the second part

OpenStudy (anonymous):

are the bounds still right? 0 to 4?

OpenStudy (anonymous):

cause i'm still getting the wrong answer. 16448/15

OpenStudy (sidsiddhartha):

whats the answer? is it 1301

OpenStudy (anonymous):

ok, 1301 is right. i ws getting 16448/15 for the integral \[\int\limits_{0}^{4}(-t^3+t^4)(1+2t)\]

OpenStudy (anonymous):

f(t)= (2t+1)⋅(t^4−t^3) ∫f(t)dt= 20t^6−12t^5−15t^4/60 ∫04f(t)dt= 16448/15 Approximation 1096.533333333333

OpenStudy (anonymous):

weird.

OpenStudy (sidsiddhartha):

no it will be \[\int\limits_{}^{}(-t^3 \hat {i}+t^4 \hat{j})*(i+2t \hat{j})\\now ~remember ~that ~ \\i*i=1\\i*j=j*i=0\\j*j=1\\so ~then~\it~will~be \\ \int\limits_{0}^{4}(-t^3+2t^5)\]

OpenStudy (anonymous):

oh, ok. i didn't know about that relationship. i'll make a note of that. thanks a bunch!

OpenStudy (sidsiddhartha):

yw!!